OpenStudy (anonymous):

if the poly. x^4-6x*+16x^2-25x+10 is divided by another poly. x^2-2x+k,the remainder comes out to be x+a,find k and a.

OpenStudy (anonymous):

what is '-6x*' is??

OpenStudy (anonymous):

* stands for cube

OpenStudy (anonymous):

well i think. if (x^2-2x+k)*(x+a) the highest power of its ans would be x^3 how could it have a x^4 ?

OpenStudy (anonymous):

@apoorvk

OpenStudy (apoorvk):

@AndrewNJ - that "x+a" is the remainder, not the quotient.

OpenStudy (anonymous):

ya..i miss it let me try to solve it

OpenStudy (anonymous):

the remainder is x+a so we can multiply (dx^2+ex+f) to simply satisfy the power of two, three and four (x^2-2x+k)(dx^2+ex+f)=dx^4-2dx^3+dkx^2+ex^3-2ex^2+ekx+fx^2-2fx+fk=dx^4+(e-2d)x^3+(dk-2e+f)x^2+(ek-2f)x+fk contrast to x^4-6x^3+16x^2-25x+10 so we can have d=1 (e-2d)=-6 (dk-2e+f)=16 (ek-2f)=-25-1=-26 fk=10-a try to solve d,e,f,k,a so, d=1 e=-4 f=3 k=5 a=-5

OpenStudy (anonymous):

thxxxxxxxxxxxxxxxxxxxxx

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i wish it has no mistake

OpenStudy (anonymous):

u r osum dude

OpenStudy (anonymous):

Isn't there a better way to do it?

OpenStudy (anonymous):

I hate it when I get too many constants.

OpenStudy (anonymous):

ya.. i dont have a better one it seems have to satisfy every coefficient of each power,it need hard work

OpenStudy (anonymous):

dis is correct

OpenStudy (anonymous):

the method is no prob but im afraid it has calculus error

OpenStudy (anonymous):

id gonna sleep. see u tomorrow, if it exists some problems :)